OpenStudy (anonymous):

What is the value of x/y -1 if 4^x=8^y. . .? Plz help.

OpenStudy (lgbasallote):

4^x = (2^2)^x 8^y = (2^3)^y 2^2x = 2^3y 2x = 3y via rule of same base exponential properties x = 3y/2 substitute... 3y/2/y - 1 3/2 - 1 = 1/2

OpenStudy (anonymous):

\[\LARGE 4^x=8^y\quad \Longrightarrow \log(4)^x=\log(8)^y\] \[\LARGE x\cdot \log(4)=y\cdot \log(8) \] \[\LARGE \frac xy =\frac{\log(8)}{\log(4)} \] \[\LARGE \frac xy =\frac{\log(2)^3}{\log(2)^2} \] \[\LARGE \frac xy =\log_{2^2}(2)^3 \] \[\LARGE \frac xy =\frac12 \cdot 3\cdot \log_{2}(2) \] \[\LARGE \frac xy =\frac32 \cdot 1 \] ... Now? :)

OpenStudy (lgbasallote):

\(\large \mathsf{\text{interesting method kreshnik /:)}}\)

OpenStudy (anonymous):

I love "log" I hate "ln" LOL

OpenStudy (lgbasallote):

lol you couldve just used the same base technique :P

OpenStudy (anonymous):

I know, anyway Logs are better when there's no way to get same bases :F (Actually I didn't even check for it, I just used Log immediately !) LOL

OpenStudy (lgbasallote):

lol =)) you think too complex :P